OpenStudy (javsteratl):

Find an equation for the line that passes through (-4, 8) and (3, -7). What is the slope? Where does the line intersect the x-axis and the y-axis?

OpenStudy (will.h):

1st you have to find the slope by the following formula \[\frac{ y2-y1 }{ x2-x1 }\] and then subtitue the values (-4,8) (3,-7) Slope= \[\frac{ -7-8 }{ 3-(-4) }\] Slope= \[\frac{ -15 }{ 7 }\]

OpenStudy (will.h):

we can find the entire function after that we found the slope, do you know how?